Domingo Corporation makes a variety of headphones with logos. The company has discovered a new market for wireless headphones with logos. Market research indicates that these headphones would sell well in the market priced at $40.00 each. Domingo desires an operating profit of 25 percent of costs.

Required:

What is the highest acceptable manufacturing cost for which Domingo would be willing to produce the headphones?

2.  


Mel’s Meals 2 Go purchases cookies that it includes in the 10,000 box lunches it prepares and sells annually. Mel’s kitchen and adjoining meeting room operate at 70 percent of capacity. Mel’s purchases the cookies for $0.72 each but is considering making them instead. Mel’s can bake each cookie for $0.23 for materials, $0.18 for direct labor, and $0.51 for overhead without increasing its capacity. The $0.51 for overhead includes an allocation of $0.33 per cookie for fixed overhead. However, total fixed overhead for the company would not increase if Mel’s makes the cookies.

Mel himself has come to you for advice. “It would cost me $0.92 to make the cookies, but only $0.72 to buy. Should I continue buying them?” Materials and labor are variable costs, but variable overhead would be only $0.18 per cookie. Two cookies are put into every lunch.

Mel suddenly finds an opportunity to sell boxed dinners. The new opportunity would require the use of the 30 percent unused capacity. The contribution margin from the dinners would amount to $3,300 annually.

Required:

a. If Mel decides to sell dinners, what are the total costs for both making and buying the cookies?
Answer #1

Solution 1

Assume that 'a' is the cost price

It is given that Domingo desires an operating profit of 25% of cost and that the headphones would sell well in the market for $40 each.

Therefore Selling price = $ 40

Operating profit = 25% *a

Selling Price = Manufacturing cost + Operating profit

$40 = a +( 25/100) * a

$40 = a+ (1/4 ) * a

$40 = a+ a/4

$40 = (4a+a)/4

$40 = 5a/4

On transposing 4 to the other side of the equation,

$ 40* 4 = 5a

$ 160 = 5a

On transposing 5 to the other side of the equation

$ 160/ 5 = a

Therefore a = $ 32

The highest acceptable manufacturing cost at which Domingo would be willing to produce the headphones would be $ 32.